i need help with this one too: If (x-2) is a factor of 3x^3 -kx^2 +(k+1)x-2; find the value of k.
Answer by ankor@dixie-net.com(12706)

3x^3 -kx^2 +(k+1)x - 2, (x-2) is one factor
:
Use synthetic division
:
..._________________
2|3 - k + (k+1) - 2
.....+ 6+ 2(6-k) +2(-k+13)
--------------------------
...3 + (6-k) + (-k+13) + 0
:
In order for us to have 0 remainder
2(-k+13) = +2
-2k + 26 = + 2
-2k = 2 - 26
-2k = -24
k = +12
:
The equation 3x^2 - 12x^2 + 13x - 2